OpenStudy (anonymous):

A golf ball is hit from the ground with an initial speed of 192 feet per second. Assume the starting height of the ball is 0 feet. How long will it take the golf ball to hit the ground?

OpenStudy (cwrw238):

use an equation of motion where acceleration is constant here: u = initial velocity = 192, a = acceleration due to gravity = -32 ft/sec^-2. s = displacement from origin = 0 and t = ? use s = ut + 0.5 a t^2 0 = 192t - 0.5*32* t^2 16t^2 - 192t = 0 16t(t - 12) = 0 t = 0 or t = 12 seconds t=0 corresponds to start so time in the air = 12 seconds